OpenStudy (anonymous):

2x2 + 12x = 54 solve by complete the square.

OpenStudy (anonymous):

*2^2

OpenStudy (anonymous):

\[x ^{2} + 6x = 27\] \[x ^{2} + 6x + 9 = 36\] \[(x + 3)(x + 3) = 36\] \[(x + 3) ^{2} = 36\] \[x + 3 = 6\] \[x = 3\]

OpenStudy (anonymous):

you may want divide everything by 2 first

OpenStudy (anonymous):

x + 3 = + or - 6

OpenStudy (wolf1728):

The method to complete the square: Move the 'non-x' term to the right Divide the equation by coefficient of x² Finally, take the coefficient of x, divide it by 2, then square it. Add this number to both sides of the equation. 2x2 + 12x = 54 x² + 6x = 27 6 divided by 2 then squared = 9 Add 9 to both sides of the equation. That should give you a good start.

OpenStudy (anonymous):

but write it as 3^2 on the left... it will remind you (hopefully) of what the factor is... (x+3)^2